OpenStudy (anonymous):

If g(x) = 3x + 5, what is g-1(x)? (1 points) 4. Select the description of f(x) = -3 - x6. (1 points) A) An odd function with end behavior of the graph that increases to the right and to the left. B) An odd function with end behavior of the graph that decreases to the right and to the left. C) An even function with end behavior of the graph that increases to the right and to the left. D) An even function with end behavior of the graph that decreases to the right and to the left. 5. What is the quotient: (3x2 - 33x + 84) ÷ (x - 7) ? (1 points)

OpenStudy (anonymous):

g-(x)=(x-5)/3 answer to the last question is 3x-12 in ur second ques is it 6x or x^6?

OpenStudy (anonymous):

x^6

OpenStudy (anonymous):

on that case option d should be correct

OpenStudy (anonymous):

thanks sooo much...and i have quite a few more questions if ud be willing to help still?

OpenStudy (anonymous):

yeah,sure

OpenStudy (anonymous):

What is the remainder when (2x3 - 3x2 - 10x + 15) ÷ (x + 2) ? What are the zeros of the polynomial function: f(x) = x3 + x2 - 6x ? What polynomial has roots of -1, -2, and 3 ? A hotel room has a volume of x3 - 5x2 - 24x - 20 ft3 and the height is (x + 2) ft. The floor is represented by what polynomial?----> try these and once again thank you sooo much

OpenStudy (anonymous):

answers are given in order of questions 7 0,-2,3 x^3-7 x^2-7x-10

OpenStudy (anonymous):

thanks omg your amazinggg lol...can u do more or no ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey post questions if u have any

OpenStudy (anonymous):

What is the simplified form of x^4/y^3/x^5/y^7? 15. What is the simplified form of 1/y+1/x over 1/y-1/x?

OpenStudy (anonymous):

answer to q1 is y^4/x

OpenStudy (anonymous):

12x over x^2+6x+9 over 4x^2 over x^2-9....simplify

OpenStudy (anonymous):

is the question

OpenStudy (anonymous):

What is the solution to the rational equation 3x over 4 - 1/4 = 5

OpenStudy (anonymous):